# Explainability of Machine Learning Models under Missing Data

### v1

@@ Abstract
Missing data is a prevalent issue that can significantly impair model performance and interpretability. This paper briefly summarizes the development of the field of missing data with respect to Explainable Artificial Intelligence and experimentally investigates the effects of various imputation methods on the calculation of Shapley values, a popular technique for interpreting complex machine learning models. We compare different imputation strategies and assess their impact on feature importance and interaction as determined by Shapley values. Moreover, we also theoretically analyze the effects of missing values on Shapley values. Importantly, our findings reveal that the choice of imputation method can introduce biases that could lead to changes in the Shapley values, thereby affecting the interpretability of the model. Moreover, and that a lower test prediction mean square error (MSE) may not imply a lower MSE in Shapley values and vice versa. Also, while Xgboost is a method that could handle missing data directly, using Xgboost directly on missing data can seriously affect interpretability compared to imputing the data before training Xgboost. This study provides a comprehensive evaluation of imputation methods in the context of model interpretation, offering practical guidance for selecting appropriate techniques based on dataset characteristics and analysis objectives. The results underscore the importance of considering imputation effects to ensure robust and reliable insights from machine learning models.
@@
Missing Data , Imputation , Shapley Values
@@
[label2]organization=Oslo Metropolitan University, city=Oslo, country=Norway
@@ 1 Introduction
Various imputation methods exist, ranging from simple techniques like mean imputation to more sophisticated approaches such as multiple imputation by chained equations (MICE) and k-nearest neighbors (KNN). Each method has its strengths and weaknesses, influencing not only the performance of predictive models but also their interpretability. For example, in [1], the authors study the effects of missing data on the correlation heatmap and found out that the technique that computes a correlation matrix with the highest accuracy (in terms of RMSE) does not always produce correlation plots that closely resemble those derived from complete data. However, the effects of missing data on various aspects of explainable AI have not yet been fully studied.
@@
Shapley values, a concept derived from cooperative game theory, have gained prominence as a robust method for interpreting complex models [2]. Shapley values attribute the contribution of each feature to the final prediction, offering insights into feature importance and interaction. However, the accuracy and reliability of Shapley values can be affected by the choice of the imputation method, as the imputed values can introduce biases or distortions. Despite the importance of this issue, little attention has been paid to the effect of imputation on the Shapley values of the downstream model. For example, in [3], the authors examine various imputation methods to choose the best one and then use Shapley values to explain the prediction of the downstream task. However, the effects of imputation methods on Shapley values have not been thoroughly examined.
@@
This paper aims to explore the effects of various imputation methods on the calculation of Shapley values. By comparing different imputation strategies, we seek to understand how they influence the interpretability of machine learning models and the robustness of the insights derived from Shapley’s values. The study provides a comprehensive evaluation of imputation methods in the context of model interpretation, offering guidance for practitioners on selecting appropriate techniques for their specific datasets and objectives.
@@
The rest of the paper is organized as follows. In Section 2, we briefly review works on explainable AI and missing data to gain insights into the development of missing data in relation to explainable AI and highlight open issues for future consideration. Next, in Section 3, we describe the methods being examined in this work. Following that, we theoretically analyze the effects of missing data on Shapley values in Section 4 and conduct and analyze the experiment results in Section 5. Finally, the paper concludes with a discussion of our findings and suggestions for future work in Section 6 and Section 7.
2 Missing data techniques and explainability
@@
In this section, we review various missing data handling techniques and their explainability.
@@ 2.1 Explainable AI
A significant development in XAI is the emergence of intrinsic interpretability approaches. These involve designing models that are interpretable by nature. Decision trees [4], rule-based models [5], and linear models [6] are classical examples of intrinsically interpretable models. More recently, models such as Generalized Additive Models (GAMs) [7] and attention mechanisms in neural networks [8] have been explored for their ability to provide insights into the decision-making process. However, their explanation follows their specific explanation style, making it difficult to compare with other models that do not follow the same explanation schemes during model selection.
@@
One prominent approach in XAI is the use of post-hoc explanation methods. These techniques, such as LIME (Local Interpretable Model-agnostic Explanations) [11] and SHAP (SHapley Additive exPlanations) [9], offer explanations by approximating the behavior of complex models locally around a given prediction. LIME creates local surrogate models that are interpretable, while SHAP leverages Shapley values from cooperative game theory to fairly distribute the contribution of each feature to the prediction. Both methods have been widely adopted due to their model-agnostic nature, allowing them to be applied to any black-box model.
Overall, compared to other techniques, the Shapley value offers several distinct advantages over other Explainable AI (XAI) techniques. Unlike methods that provide only local explanations or are limited to specific model architectures, the Shapley value delivers a unified and theoretically grounded approach applicable to any machine learning model. It ensures fairness by distributing the model’s output among the features based on their marginal contributions, offering a comprehensive and additive feature attribution that respects interaction effects. This fairness and consistency make Shapley values particularly valuable in high-stakes domains like medicine or finance [12, 13, 14, 15, 16, 17], where understanding the relative importance and interplay of various features is crucial for trust and accountability. Additionally, Shapley values provide a global interpretability framework, making it easier to understand the model’s behavior across different scenarios, thus enhancing transparency and facilitating more informed decision-making.
However, the robustness and reliability of these attributions can be significantly influenced by the treatment of missing data.
@@ 2.2 Missing data imputation techniques
The most common approach to handling missing values is to use imputation methods to fill in the gaps. Techniques such as matrix decomposition or matrix completion, including MGP [18], Polynomial Matrix Completion [19], ALS [20], and Nuclear Norm Minimization [21], allow for continuous data to be completed and subsequently analyzed using standard data analysis procedures. Additionally, regression or clustering-based methods, such as CBRL and CBRC [22], utilize Bayesian Ridge Regression and cluster-based local least square methods [23] for imputation. For large datasets, deep learning imputation techniques have gained popularity due to their performance [24, 25, 26]. It’s important to note that different imputation methods may produce varying values for each missing entry, making the modeling of uncertainty for each missing value crucial. Bayesian and multiple imputation techniques, such as Bayesian principal component analysis-based imputation [27] and multiple imputations using Deep Denoising Autoencoders [28], are particularly useful in these scenarios. Furthermore, certain tree-based methods can naturally handle missing data through prediction, including missForest [29], the DMI algorithm [30], and DIFC [31]. Methods that can manage mixed data types have also emerged, such as SICE [32], FEMI [33], and HCMM-LD [34].
@@
Although various missing data methods have been developed [35], most of them to date do not have a built-in explanation mechanism for imputed values, although how a value is imputed can have a profound impact on the performance and interpretability of downstream machine learning models. If one uses a prediction model such as a regression or classification model, then one can use Shapley values or some other XAI technique to explain the imputation of missing values. However, many times, one may want to use a more complicated imputation method that requires building regression or classification models and looping through the data multiple times to improve imputation accuracy.
@@
Sometimes, the imputed values of some imputation techniques can be easily achieved even though the original algorithm does not provide an explanation. K-nearest neighbor imputation may offer some explanation of the imputed values based on the K-nearest neighbors. Next, since tree-based techniques are explainable, the explanation for tree-based imputation methods such as missForest [29] can be achieved by building a tree to explain the classification results. However, a potential ambiguity can be the prediction of one missing entry based on the imputed values of the other entries during the iteration process. Similar things apply to MICE imputation [36], a method for handling missing data through iterative predictions based on the relationships between observed and missing variables.
In recent years, more attention has been paid to the explainability of the imputation method under missing data. For example, [37] introduces DIMV, an imputation method that provides explainability and uncertainty of imputation via conditional distribution, or [1] analyzes the effects of various imputation and direct parameter estimation methods to the correlation plot. Next, Hans et al. [38] introduce an explainable imputation method that leverages association constraints within the data. However, up to our knowledge, so far, there is no work on the effects of missing data on Shapley values.
@@ 2.3 Direct missing data handling techniques without imputation
Different from the imputation approaches, methods that directly handle missing data can have clearer implications in terms of explainability.
Specifically, Nguyen et al. [35] introduced the EPEM algorithm to estimate the maximum likelihood estimates (MLEs) for multiple class monotone missing data when the covariance matrices of all classes are assumed to be equal. Additionally, DPER [39] addresses a more general case where missing data can occur in any feature by using pairs of features to estimate the entries in the mean and covariance matrices. The implication to model explainability of using direct parameter estimation methods, like EPEM and DPER, includes improved transparency and interpretability of the model’s behavior, as these methods provide clear estimates of means and covariances, which can be directly examined and understood. In fact, [1] analyzes the effects of various imputation and direct parameter estimation methods on the correlation plot, and finds that DPER is not only scalable but also better for this task than imputation techniques such as SOFT-IMPUTE [40], KNNI, GAIN [41], GINN [42], missForest [29].
Note that LSTM [43] can handle missing data directly. In fact, it is used the work of Ghazi et al. [44] to model disease progression while handling missing data in both the inputs and targets. Another work by Li et al. [45] proposed a technique for the bi-clustering problem that can manage missing data, enabling simultaneous partitioning of rows and columns of a rectangular data array into homogeneous subsets. Learning directly from the data may offer advantages in speed and reduced storage costs by eliminating the need for separate models for imputation and the target task.
@@ 3.1 Shapley values
Shapley values, originally derived from cooperative game theory [46], have been adapted as a powerful tool for interpreting machine learning models [9]. They provide a systematic way to attribute the contribution of each feature to a model’s prediction, ensuring a fair distribution based on the interaction of features. Consider a machine learning model $v$ that takes an input vector $\mathbf{x}=(x_{1},x_{2},...,x_{p})$ and produces a prediction $v(\mathbf{x})$. The Shapley value for the value $x_{i}$ of corresponding feature $\boldsymbol{f}_{i}$ quantifies its contribution to the prediction $v(\mathbf{x})$. Formally, let $P=\{1,2,\dots,p\}$ be the complete index set of features and $S$ a coalition of feature indices (a subset of $P\setminus\{i\}$). For $i=1,2,\dots,p$, the Shapley value $\phi_{i}$ for the feature value $x_{i}$ is defined as
$\phi_{i}=\sum_{S\subseteq P\setminus\{i\}}\frac{|S|!(p-|S|-1)!}{p!}\left[v(S\cup\{i\})-v(S)\right],$
where $|S|$ is the number of elements in subset $S$, and $v(S\cup\{i\})$, $v(S)$ is the model prediction using feature indices in $S$ plus $i^{th}$ feature and only in $S$, respectively. In the context of classification problems, the model outputs a probability distribution over classes. The Shapley values can be calculated for each class’s probability, providing insights into how each feature influences the likelihood of each class. Specifically, calculating Shapley values in classification problems can be broken down into the algorithm 1.
@@
    A Global Feature Importance plot is a visualization tool for examining the impact of individual features within a predictive model. More specifically, it highlights the features that contribute the most to the model’s predictions, enabling a deeper understanding of the model’s behavior. Here, the global importance of each feature is determined by calculating the mean absolute value of Shapley values of that feature across all the given samples.
@@
    A beeswarm plot is a useful visualization tool to depict the distribution and influence of individual feature contributions on the model’s predictions, particularly when using Shapley values. It provides a detailed view of how each feature affects the output of a model across the entire dataset. By displaying each data point as a dot and arranging these dots to show the distribution of Shapley values for each feature, the beeswarm plot offers a comprehensive overview of feature importance and interaction effects. It allows for the identification of patterns and outliers, helping to understand the behavior of the model with respect to individual features. Additionally, the plot can reveal how specific features interact with others, providing insights into complex dependencies within the data.

### v2

@@ Abstract
Missing data is a prevalent issue that can significantly impair model performance and interpretability. This paper briefly summarizes the development of the field of missing data with respect to Explainable Artificial Intelligence and experimentally investigates the effects of various imputation methods on SHAP (SHapley Additive exPlanations), a popular technique for explaining the output of complex machine learning models. Next, we compare different imputation strategies and assess their impact on feature importance and interaction as determined by Shapley values. Moreover, we also theoretically analyze the effects of missing values on Shapley values. Importantly, our findings reveal that the choice of imputation method can introduce biases that could lead to changes in the Shapley values, thereby affecting the explainability of the model. Moreover, we also show that a lower test prediction MSE (Mean Square Error) does not necessarily imply a lower MSE in Shapley values and vice versa. Also, while XGBoost (eXtreme Gradient Boosting) is a method that could handle missing data directly, using XGBoost directly on missing data can seriously affect interpretability compared to imputing the data before training XGBoost. This study provides a comprehensive evaluation of imputation methods in the context of model explanations, offering practical guidance for selecting appropriate techniques based on dataset characteristics and analysis objectives. The results underscore the importance of considering imputation effects to ensure robust and reliable insights from machine learning models.
@@
Missing Data , Imputation , Model Explainability
@@
[…]
@@ 1 Introduction
Various imputation methods exist, ranging from simple techniques like mean imputation to more sophisticated approaches such as multiple imputation by chained equations (MICE) and k-nearest neighbors (KNN). Each method has its strengths and weaknesses, influencing not only the performance of predictive models but also their interpretability. For example, in [1], the authors studied the effects of missing data on the correlation heatmap and found out that the technique that computes a correlation matrix with the highest accuracy (in terms of root mean square error, RMSE) does not always produce correlation plots that closely resemble those derived from complete data. However, the effects of missing data on various aspects of explainable AI (XAI) have not yet been fully studied.
@@
In terms of explainability, Shapley values, a concept derived from cooperative game theory, have gained prominence as a robust method for interpreting complex models [2]. Shapley values attribute the contribution of each feature to the final prediction, offering insights into feature importance and interaction. However, the accuracy and reliability of Shapley values can be affected by the choice of the imputation method, as the imputed values can introduce biases or distortions. Despite the importance of this issue, little attention has been paid to the effect of imputation on the Shapley values of the downstream model. For example, in [3], the authors examine various imputation methods to choose the best one and then use Shapley values to explain the prediction of the downstream task. However, the effects of imputation methods on Shapley values have not been thoroughly examined.
@@
This paper aims to explore the effects of various imputation methods on the calculation of Shapley values of the downstream classification/regression task. By comparing different imputation strategies, we seek to understand how they influence the interpretability of machine learning models and the robustness of the insights derived from Shapley values. The study provides a comprehensive evaluation of imputation methods in the context of model interpretation, offering guidance for practitioners on selecting appropriate techniques for their specific datasets and objectives.
In summary, the contributions of this paper are: (i) we explore the effects of various imputation methods on Shapley values of the downstream regression/classification task; (ii) we demonstrate that the choice of imputation method can significantly impact Shapley values, influencing the interpretability of the downstream machine learning models; (iii) we highlight that imputation methods achieving the highest prediction accuracy do not necessarily produce the most reliable feature importance values, suggesting the need for method selection aligned with dataset characteristics and analysis goals.
@@
The rest of the paper is organized as follows. In Section 2, we review related works on XAI and missing data to gain insights into the development of missing data in relation to XAI and highlight open issues for future consideration. Next, in Section 3, we describe the methods being examined in this work. Following that, we theoretically analyze the effects of missing data on Shapley values in Section 4 and conduct and analyze the experiment results in Section 5. Finally, the paper concludes with a discussion of our findings and suggestions for future work in Section 6 and Section 7.
2 Related works
@@
In this section, we review various related works on missing data handling techniques and their explainability.
@@ 2.1 Explainable AI
A significant development in XAI is the emergence of intrinsic interpretability approaches, which involve designing models that are interpretable by nature. Decision trees [4], rule-based models [5], and linear models [6] are classical examples of intrinsically interpretable models. More recently, models such as Generalized Additive Models (GAMs) [7] and attention mechanisms in neural networks (NNs) [8] have been explored for their ability to provide insights into the decision-making process. However, their explanation follows their specific explanation style, making it difficult to compare with other models that do not follow the same explanation schemes during model selection.
@@
One prominent approach in XAI is the use of post-hoc explanation methods, such as LIME (Local Interpretable Model-agnostic Explanations) [11] and SHAP (SHapley Additive exPlanations) [9], offer explanations by approximating the behavior of complex models locally around a given prediction. LIME creates interpretable local linear surrogate models, while SHAP leverages Shapley values from cooperative game theory to fairly distribute the contribution of each feature to the prediction. Both methods have been widely adopted due to their model-agnostic nature, allowing them to be applied to any black-box model.
Overall, compared to other techniques, the Shapley value offers several distinct advantages over other Explainable AI (XAI) techniques. Unlike methods that provide only local explanations or are limited to specific model architectures, the Shapley value delivers a unified and theoretically grounded approach applicable to any machine learning model. It ensures a mathematical notion of fairness by distributing the model’s output among the features based on their marginal contributions over different patterns of availability of the other features, offering an additive feature attribution that respects interaction effects. This fairness and consistency make Shapley values particularly valuable in high-stakes domains like medicine or finance [12, 13, 14, 15, 16, 17], where understanding the relative importance and interplay of various features is crucial for trust and accountability. Additionally, Shapley values provide a global interpretability framework, making it easier to understand the model’s behavior across different scenarios, thus enhancing transparency and facilitating more informed decision-making. However, the treatment of missing data can significantly influence the robustness and reliability of XAI techniques, including Shapley values.
@@ 2.2 Missing data imputation techniques
The most common approach to handling missing values is to use imputation methods to fill in the gaps. Techniques such as matrix decomposition or matrix completion, including missing Gaussian processes (MGP) [18], Polynomial Matrix Completion [19], matrix completion via alternating least squares (ALS) [20], and a congestion imputation model (CIM) based on joint matrix factorization [21], allow for continuous data to be completed and subsequently analyzed using standard data analysis procedures. Additionally, regression or clustering-based methods, such as cumulative Bayesian ridge with less NaN (CBRL) and cumulative Bayesian ridge with high correlation (CBRC) [22], utilize Bayesian Ridge Regression and cluster-based local least square methods [23] for imputation. For large datasets, deep learning imputation techniques have gained popularity due to their performance [24, 25, 26]. It is important to note that different imputation methods may produce varying values for each missing entry, making the modeling of uncertainty for each missing value crucial. Bayesian and multiple imputation techniques, such as Multiple Imputation with Denoising Autoencoders (MIDAS) [25], Bayesian principal component analysis-based imputation [27] and multiple imputations using Deep Denoising Autoencoders [28], are particularly useful in these scenarios. Furthermore, certain tree-based methods can naturally handle missing data through prediction, including missForest [29], the decision tree missing-value imputation (DMI) algorithm [30], and decision trees and iterative fuzzy clustering (DIFC) [31]. Methods that can manage mixed data types have also emerged, such as Single Center Imputation from Multiple Chained Equation (SICE) [32], and Clustering Mixed Numerical and Categorical Data with Missing Values (k-CMM) [33].
@@
Although various missing data methods have been developed, most of them do not have a built-in explanation mechanism for imputed values. However, the way a value is imputed can have a profound impact on the performance and interpretability of downstream machine learning models. If one uses a prediction model such as a regression or classification model, then one can use Shapley values or some other XAI technique to predictions obtained from models trained after imputing training data. However, many times, one may want to use a more complicated imputation method that requires building regression or classification models and looping through the data multiple times to improve imputation accuracy. This can make it more challenging to explain the imputed values than imputing the data without using loops such as based on a regression model, for example.
@@
Sometimes, the explanation for imputed values of some imputation techniques can be easily achieved even though the original algorithm does not provide an explanation. K-nearest-neighbor imputation may offer some explanation of the imputed values based on the $K$ nearest neighbors. Next, since tree-based techniques are explainable, the explanation for tree-based imputation methods such as missForest [29] can be achieved by building a tree to explain the classification results. However, a potential ambiguity can be the prediction of one missing entry based on the imputed values of the other entries during the iteration process. Similar concerns apply to MICE imputation [34], a method for handling missing data through iterative predictions based on the relationships between observed and missing variables.
Attention to the explainability of imputation techniques under missing data has intensified in recent years. For example, [35] introduces Distribution-based Imputation of Missing Values with Regularization (DIMV), an interpretable imputation method that characterizes the uncertainty of imputation based on the conditional distribution of a multivariate Gaussian model, with regularization and feature selection; or [1] analyzes the effects of various imputation and direct parameter-estimation methods to the correlation plot. Next, a technique for obtaining local explanations when the point to be explained and/or the reference dataset are affected by data missingness is proposed in [36]. More specifically, they propose an extension of LIME that can be used to explain the output of models that handle missing data without imputation, such as the ones introduced in the ensuing section.
However, to our knowledge, so far, there has not been much work on the effects of missing data on Shapley values. In fact, while [37] implements five imputation methods (elimination, mean, mode, KNN, MICE) to deal with missing data, they made predictions using various prediction models to conclude that XGBoost combined with KNN imputation gives the best results; they only examine SHAP for this combination.
@@ 2.3 Direct missing data handling techniques without imputation
Different from the imputation approaches, methods that directly handle missing data can have clearer implications in terms of explainability. Specifically, Nguyen et al. [38] introduced an efficient Parameter Estimation for Multiple Class Monotone Missing Data (EPEM) algorithm to estimate the maximum likelihood estimates (MLEs) for multiple class monotone missing data when the covariance matrices of all classes are assumed to be equal. Additionally, Direct Parameter Estimation for Randomly missing data (DPER) [39] addresses a more general case where missing data can occur in any feature by using pairs of features to estimate the entries in the mean and covariance matrices. The implication to model explainability of using direct parameter estimation methods, like EPEM and DPER, includes improved transparency and interpretability of the model’s behavior, as these methods provide clear estimates of means and covariances, which can be directly examined and understood. In fact, [1] analyzes the effects of various imputation and direct parameter estimation methods on the correlation plot, and finds that DPER is not only scalable but also better for this task than imputation techniques such as SOFT-IMPUTE [40], KNN-based imputation (KNNI), generative adversarial imputation network (GAIN) [41], graph imputer neural network (GINN) [42], and missForest [29].
Note that long short-term memory (LSTM) networks [43] can handle missing data directly. In fact, it is used in the work of Ghazi et al. [44] to model disease progression while handling missing data in both the inputs and targets. The bi-clustering problem [45], a statistical learning methodology that enables simultaneous partitioning of rows and columns of a rectangular data array into homogeneous subsets, is tackled by Li et al. [46] while handling missing data. Learning directly from the data may offer advantages in speed and reduced storage costs by eliminating the need for separate models for imputation and the target task.
@@
2.4 Studies on the impact of imputation on model interpretation
In the work of Shadbahr et al., it is observed that poor imputation quality can lead to misleading feature importance assessments in ML classifiers [47]. More specifically, they design a simulated dataset where they expect the distribution of the Shapley values to be symmetric, as the clusters are separated by a fixed distance and are normally distributed within those clusters. Using the skewness, they measure the symmetry of the feature importance values for each feature for all models. According to their experiments, models fit poorly imputed data giving rise to misleading feature importance metrics, demonstrating the adverse impact of poor imputation quality on model interpretability.
As highlighted by Payrovnaziri et al., changes in the underlying data due to imputation can alter model characteristics and feature importance, necessitating a careful examination of how imputed values influence predictions [48]. Their results show that the level of missingness and the imputation method used can have a significant impact on the interpretation of the models. To assess model interpretability, they focused on the stability of feature importance rankings across multiple imputations. Their findings revealed that the choice of imputation method significantly influences the interpretability of predictive models, with MICE providing more stable feature importance rankings compared to mean imputation and k-NN.
Assessing the impact of imputation in the model explainability is particularly relevant in clinical settings, where accurate interpretation of model outputs is essential for decision-making. The effectiveness of interpretable ML models under missing data was also studied in [49] from a subjective point of view (surveying clinicians’ perception).
The impact of missing data imputation in the performance and interpretability of prediction rule ensembles (PREs) was studied in [50]. A simulation study is conducted where a PRE model is trained using data imputed using several imputation schemes, and the model size and performance are evaluated, observing a trade-off between performance and model complexity.
@@ 3.1 Shapley values
Shapley values, originally derived from cooperative game theory [51], have been adapted as a powerful tool for interpreting machine learning models [9]. They provide a systematic way to attribute the contribution of each feature to a model’s prediction, in accordance with a mathematical notion of fairness in the distribution based on the interaction of features. Consider a machine learning model $v$ that takes an input vector $\mathbf{x}=(x_{1},x_{2},...,x_{p})$ and produces a prediction $v(\mathbf{x})$. The Shapley value for the value $x_{i}$ of corresponding feature $\boldsymbol{f}_{i}$ quantifies its contribution to the prediction $v(\mathbf{x})$. Formally, let $P=\{1,2,\dots,p\}$ be the complete index set of features and $S$ a coalition of feature indices (a subset of $P\setminus\{i\}$). For $i=1,2,\dots,p$, the Shapley value $\phi_{i}$ for the feature value $x_{i}$ is defined as
$\phi_{i}=\sum_{S\subseteq P\setminus\{i\}}\frac{|S|!(p-|S|-1)!}{p!}\left[v(S\cup\{i\})-v(S)\right],$ (1)
where $|S|$ is the number of elements in subset $S$, and $v(S\cup\{i\})$, $v(S)$ denote the model prediction using feature indices in $S$ plus $i^{th}$ feature and only in $S$, respectively. In the context of classification problems, the model outputs a probability distribution over classes. The Shapley values can be calculated for each class’s probability, providing insights into how each feature influences the likelihood of each class. Specifically, calculating Shapley values in classification problems can be broken down into Algorithm 1.
@@
    A global feature importance plot allows visualizing the impact of individual features within a predictive model. More specifically, it highlights the features that contribute the most to the model’s predictions, enabling a deeper understanding of the model’s behavior. Here, the global importance of each feature is determined by calculating the mean absolute value of Shapley values of that feature across all the given samples.
@@
    A beeswarm plot is a useful visualization tool to depict the distribution and influence of individual feature contributions on the model’s predictions, particularly when using Shapley values. It provides a detailed view of how each feature affects the output of a model across the entire dataset. By displaying each data point as a dot and arranging these dots to show the distribution of Shapley values for each feature, the beeswarm plot offers a comprehensive overview of feature importance. It allows for the identification of patterns and outliers, helping to understand the behavior of the model with respect to individual features.